OpenStudy (anonymous):

Determine the equation of the secant line through the given x values: f(x)=3x^2-5x+4; x=-3, x=2

OpenStudy (anonymous):

Would I just find the slope, then plug in one value of x into f(x), then create an equation using point slope form?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no worries. make sure whatever value of x you pick in your point-slope form you use the corresponding value of y=f(x)

OpenStudy (anonymous):

\[y-f(x_1)=\frac{f(x_2)-f(x_1)}{2-(-3)}(x-x_1)\]